# Out of spuriousity: Improving robustness to spurious correlations without group annotations

### v1

@@ 4 Background: Subnetwork Extraction
In this section, we outline the approach for the extraction of subnetworks (Csordás et al., 2020). Generally, we aim to extract a sparsely connected subnetwork from a dense network with a binary mask that indicates which weights to include (the subnetwork) and which to discard.
@@
Given a trained neural network $f$, for layer $l$ in $L$ hidden layers of a neural network $f$, we introduce a learnable parameter $\bm{\Pi=(\pi^{1},\pi^{2},...,\pi^{L})}$ corresponding to each layer’s weights $\mathbf{W=(w^{1},w^{2},...,w^{L})}$. Each element of $\bm{\pi^{l}}$ is assumed to be an independent Bernoulli random variable. Each element $\pi_{j,k}^{l}$ performs as a logit which indicates the probability of keeping the corresponding weight $w_{j,k}^{l}$. Weight $w_{j,k}^{l}$ connects the $j^{th}-$neuron from $(l-1)^{th}-$layer to $k^{th}-$neuron from $l^{th}-$layer, we write each weight and its corresponding logit in short $w_{i}$ and $\pi_{i}$. During training, network parameters in hidden layers become $(\mathbf{\bm{w^{1}\odot\pi^{1},w^{2}\odot\pi^{2},...,w^{L}\odot\pi^{L}}})$.
Mask training.
To train the binary mask, we freeze the model weights $\mathbf{W=(w^{1},w^{2},...,w^{L})}$ and train only the added parameters $\bm{\Pi=(\pi^{1},\pi^{2},...,\pi^{L})}$ with each $\pi_{i}$ equal to 0.9 initially, i.e., high keep probability. Modular loss $\mathrm{L_{mod}}$ (Csordás et al., 2020) aims to extract a set of sparse weights (the subnetwork) that retain the performance of the original densen netwrok on the classification task.
$\mathrm{L_{mod}}=\mathrm{L_{CE}}+\alpha\sum_{i}\pi_{i},$
where $\mathrm{L_{CE}}$ is cross-entropy loss, $\sum_{i}\pi_{i}$ sparse regularization which keeps a logit $\pi_{i}$ small unless it is needed for the task and $\alpha$ is responsible for the strength of the sparse regularization.
@@
$s_{i}=\sigma((\pi_{i}-\log(\log U_{1}/\log U_{2})/{\tau})\text{ with }U_{1},U_{2}\sim\text{U}(0,1),$
@@
$m_{i}=[\mathds{1}_{s_{i}>\gamma}-s_{i}]_{stop}+s_{i}$
where threshold $\gamma$ is 0.5 by default, $\mathds{1}$ is the indicator function and [.$[.]_{stop}$ is the stop gradient operator.
The final parameters of the resulting subnetwork are defined as $\mathbf{W^{\prime}=(w^{1}\odot m^{1},w^{2}\odot m^{2},...,w^{L}\odot m^{L})}$.
@@ 5 Our Approach
We hypothesize that instances with the same spurious attribute are nearby in representation space, and that spurious features induce clusters (cf. illustration in Fig.3 A). We verify this hypothesis for multiple attributes on the CelebA dataset (Sec. 5.1).
Our approach centers around two key ideas: First, we identify clusters in the latent space and move instances with the same spurious attributes away from each other and samples of the same class closer to each other. Second, we reduce the representation capacity of the network to a subnetwork that is less prone to spurious correlations.
More specifically, we learn a task-oriented subnetwork (Sec. 5.2) with a task-specific contrastive loss (Sec. 5.2.1) that optimizes the representation space. We detail the specific data selection and sampling procedure for negative and positive samples in the contrastive loss in Sec. 5.2.2. Finally, we present the overall training procedure in Sec. 5.3.
5.1 Analysis of ERM representation space for multipe spurious features
@@
In the presence of spurious features, ERM tends to learn these as predictive features yielding decision boundaries aligned with the spurious feature (as exemplified on the blue moons data set in Fig. 1). We analyze the ERM behavior in the presence of multiple spurious features on the CelebA dataset. We use an ERM model for predicting hair color (blond or not blond) having a training accuracy of 95%.
To analyze the feature space, we choose three other attributes (gender, age, presence of a beard) as potential spurious attributes. For a brief description of the CelebA dataset see Sec. 6.1 and Appendix A.1 detailed information. Fig. 2 shows the t-SNE (Van der Maaten & Hinton, 2008) projection of a random subset of CelebA training data representations from the penultimate layer111 We discard samples with more than one of the selected attributes for easier interpretation.. The two classes are well separated in representation space (Fig. 2, left). Interestingly, there is a strong correlation of spurious features and classes in representation space. Female correlates with blond hair, whereas the cluster for non-blond hair (orange cluster, left in Fig. 2) corresponds to young people and people with beard. Moreover, within one class (non-blond hair) the two attributes partition the data in manifolds even though attribute labels are not using during training. Lastly, unsupervised k-means clustering in representation space tends to cluster by attributes with high purity (Fig. 2, right). This means, when optimizing for class separation, instances from the same attribute group $a_{i}\in\mathcal{A}$ (e.g., persons with beard, young persons) lie close in representation space, even though the group labels are not used during training.
@@
Tab. 1 shows the purity of class (blond or non-blond hair) and spurious attributes (male or female) in clusters obtained from $k$-means clustering on representations of the CelebA training dataset. Each cluster is assigned a label based on the most frequent class (or spurious) attribute within it. Purity is calculated as the number of correctly matched attribute and cluster labels divided by the total number of instances in that cluster. The table demonstrates that $k$-means clustering effectively groups samples by class label, but achieves even higher purity for spurious attributes. Fig. 2 (right) demonstrates $k$-means clustering ($k=8$) well defines clusters based on not only one but multiple spurious attributes.
The critical case is when the model potentially can separate both, classes and attribute groups, i.e., when the model has high predictive performance and the latent space shows clear clusters for spurious features. Such a model will i) make predictions based on spurious features and ii) not generalize well to samples with attribute combinations that do not align with the spurious correlations.
@@
Based on these observations, our goal is to ensure that spurious features do not define data manifolds and cannot be used as discriminative features.
5.2 Task-oriented Subnetwork
@@
To ensure that spurious features do not define data manifolds and cannot be used as discriminative features, we introduce a task-oriented subnetwork, and a contrastive loss that optimizes the representation space. More specifically, we extend the approach for extracting subnetworks (cf. Sec. 4) by incorporating a task-specific loss function to prevent the network from learning a grouping of samples based on spurious features. The loss for our task-oriented subnetwork is defined as
$\mathrm{L}=\mathrm{L_{mod}}+\beta\mathrm{L_{task}},$
@@
where $\mathrm{L_{mod}}$ is the modular loss (cf. Sec. 4), $\mathrm{L_{task}}$ the task-specific loss, and $\beta$ is a hyper-parameter to balance loss terms.
@@
The choice of $\mathrm{L_{task}}$ in combination with a mask-training dataset $\mathcal{D}_{task}$ defines the specific task the subnetwork is optimized for. For instance, Csordás et al. (2020) find a subnetwork responsible for a specific class ($\mathcal{D}_{task}$ contains examples of this class, $\mathrm{L_{task}}$ = 0), and Park et al. (2023) identify a subnetwork that is more robust to a spurious feature ($\mathcal{D}_{task}$ contains instances misclassified by ERM, and $\mathrm{L_{task}}$ is a debiasing term). Details on those methods can be found in Appendix A.2. We detail our choice of $\mathrm{L_{task}}$ in Sec. 5.2.1 and $\mathcal{D}_{task}$ in Sec. 5.2.2.
@@
Fig. 3 illustrates the goal and training procedure of our task-oriented subnetwork. Based on the assumption that instances from the same class (same color) lie in different clusters, and that spurious features induce clusters, we define a contrastive loss $\mathrm{L_{task}}$, such that samples from the same class in different clusters are moved closer in representation space, while samples from the same cluster are moved away from each other.
@@
Figure 3: Assumption and overall idea. A. Instances from the same class (same colour) lie in different clusters apart in feature space. Instances with the spurious feature (solid border) are nearby in feature space, i.e., the spurious feature induces clusters and primarily defines the shape of the data manifold. B. Contrastive loss: For an anchor $x$, instances from the same class but different clusters constitute positive samples $x^{+}$, instances from the same cluster negative samples $x^{-}$. Positive samples are moved closer to the anchor, negative samples away from the anchor. In each iteration multiple samples act as anchors (with their positive and negative samples). C. Goal and effect of our contrastive loss. The spurious feature does not define data manifolds and cannot be used as discriminative feature between classes. Note: Our approach does not assume (and need) knowledge about the spurious feature.
@@ 5.2.1 Contrastive Learning
We use a variant of contrastive loss (Khosla et al., 2020) to mitigate the tendency of ERM to form clusters based on spurious attributes. Different from the original work, our contrastive batch is not only defined by class labels, but also by the location of samples in representation space. Fig. 3 outlines the general idea.
For each contrastive batch we randomly sample a set of anchors and for each anchor a corresponding set of positive and negative samples. We define positive samples as having the same class as the anchor, but from a different cluster than the anchor. We define negative samples as samples from the same cluster (independent of their class label). For anchor $(x,y)$ belonging to cluster $\mathrm{c}\subset\mathcal{C}$, the positive set of size P is $\left\{(x^{+},y^{+})\mid(x^{+},y^{+})\in\mathcal{C}\setminus\mathrm{c}\land y^{+}=y\right\}$ and the negative set of size N is $\left\{(x^{-},y^{-})\mid(x^{-},y^{-})\in\mathrm{c}\setminus(x,y)\right\}$. The contrastive loss term defined with respect to the anchor $(x,y)$ is
$\mathrm{l_{con}}(x)=-\sum\limits_{i=1}^{P}\log\frac{\exp(z^{\top}z_{i}^{+}/\tau)}{\sum_{P}\exp(z^{\top}z_{p}^{+}/\tau)+\sum_{N}\exp(z^{\top}z_{n}^{-}/\tau)}$
where $\tau>0$ is a scalar temperature hyperparameter, $z,z^{+},z^{-}$ are the normalized representations of corresponding inputs $x$, $x^{+}$ and $x^{-}$ from anchor, positives and negatives respectively.
This loss forces the representations of samples of the same class across clusters to become more similar and more similar to the anchor, while samples representing potential spurious attributes are pushed away from the anchor.
@@ 5.2.2 Data
PruSC does not require prior knowledge of spurious features and is applicable in the presence of multiple spurious features. We do not assume that group annotations are available but automatically infer groups to construct the de-biasing dataset $\mathcal{D}_{task}$. Based on the assumption that the representation space of ERM has regions corresponding to spurious attributes (cf. Sec. 5.1, and Fig. 2), we use unsupervised clustering to construct the spurious groups for our dataset $\mathcal{D}_{task}$.
@@
Given an ERM model, we calculate the embedding of the training data $\mathcal{D}$ and apply k-means clustering on the embedding. Each sample in the training set $\mathcal{D}$ is assigned to a cluster $c$, $(x,y,c)\in\mathcal{D}=\mathcal{X}\times\mathcal{Y}\times\mathcal{C}$, with input domain $\mathcal{X}$, class labels $\mathcal{Y}$, and clusters $\mathcal{C}$. We define $\mathcal{C}=\{c_{j}|j\in I\}$ with the cluster index set $j\in\mathrm{I}=\{1,2,\ldots,n\}$.
@@
We differentiate clusters based on their class purity. We call a cluster $\mathrm{c}$ $i$-dominant for class $i$, if at least 90% of its samples belong to class $i$, i.e., a cluster is class-dominant, if $\exists i\in\mathcal{Y}$ such that |$\frac{|\{(x,y,c)\in\mathrm{c}\mid y=i\}|}{|\mathrm{c}|}\geq 0.9$ then $\mathrm{c}$ is $i-$dominant. Otherwise, $\mathrm{c}$ is neutral. A cluster can be either neutral or $i-$dominant for some $i\in\mathcal{Y}$ and multiple different clusters may be $i-$dominant for the same class $i$. $\mathrm{C}^{i}$ is the union of $i-$dominant clusters (of the same class $i$) and $\mathrm{C}^{N}$ is the union of neutral clusters
$\mathrm{C}^{i}=\bigcup\limits_{j\in\mathrm{I}_{i}}\mathrm{c}_{j}\qquad\mathrm{C}^{N}=\bigcup\limits_{j\in\mathrm{I}_{N}}\mathrm{c}_{j},$
with $\mathrm{I}_{i}\subset\mathrm{I}$ being the set index of $i-$dominant clusters and $\mathrm{I}_{N}\subset\mathrm{I}$ being the set index of neutral clusters.
Let $m_{i}$ be the minority set of samples with label $i$ assigned to clusters dominated by a different class
$m_{i}=\{(x,y,c_{j})\mid j\not\in(\mathrm{I}_{i}\cup\mathrm{I}_{N})\land y=i\}$
@@
To construct a class-balanced dataset $\mathcal{D}_{task}$ for subnetwork training, we determine a fixed number $p$ of samples for each class. We draw $p$ samples for class $i$ as follows: First, we take all samples from the minority set $m_{i}$. Second, for each cluster in $\mathrm{C}^{i}\cup\mathrm{C}^{N}$, we sample a number of $\frac{p-|m_{i}|}{|\mathrm{I}_{i}|+|\mathrm{I}_{N}|}$ samples having class label $i$. We obtain a subset of the training data $\mathcal{D}_{task}$ which is class-balanced with $p$ samples for each class.

### v2

@@ 4 Background: Subnetwork Extraction
In this section, we outline the approach for the extraction of subnetworks (Csordás et al., 2020), i.e., subnetworks responsible for a specific task. The extraction method is also utilized by Zhang et al. (2021) to solve out-of-distribution generalization by probing subnetworks serving different functional subtasks from a pre-trained model. Additionally, Park et al. (2023) use this approach to train a subnetwork that more robust to a specific set of hard cases, therefore, mitigating the effects of spurious correlations in learning. In generally, we aim to extract a sparsely connected subnetwork from a dense network with a binary mask indicating which weights to include (the subnetwork) and which to discard.
@@
Given a trained neural network $f$, for layer $l$ in $L$ hidden layers of a neural network $f$, we introduce a learnable parameter $\bm{\Pi=(\pi^{1},\pi^{2},...,\pi^{L})}$ corresponding to the weights of each layer $\mathbf{W=(w^{1},w^{2},...,w^{L})}$. Each element $\pi_{j,k}^{l}$ acts as a logit indicating the probability of keeping the corresponding weight $w_{j,k}^{l}$. The weight $w_{j,k}^{l}$ connects the $j^{th}$-neuron from the $(l-1)^{th}$-layer to the $k^{th}$-neuron from $l^{th}$-layer, we write each weight and its corresponding logit as $w_{i}$ and $\pi_{i}$ for short. During training, network parameters in hidden layers become $(\mathbf{\bm{w^{1}\odot\pi^{1},w^{2}\odot\pi^{2},...,w^{L}\odot\pi^{L}}})$.
The Mask Training.
To train the binary mask, we freeze the model weights $\mathbf{W=(w^{1},w^{2},...,w^{L})}$ and train only the added parameters $\bm{\Pi=(\pi^{1},\pi^{2},...,\pi^{L})}$ with each $\pi_{i}$ equal to 0.9 initially, i.e., high keep probability. Modular loss $\mathrm{L_{mod}}$ (Csordás et al., 2020) aims to extract a set of sparse weights (the subnetwork) that retain the performance of the original dense network on the classification task.
$\mathrm{L_{mod}}=\mathrm{L_{CE}}+\alpha\sum_{i}\pi_{i},$ (1)
where $\mathrm{L_{CE}}$ is the cross-entropy loss, $\sum_{i}\pi_{i}$ is sparse regularization, which keeps a logit $\pi_{i}$ small unless it is needed for the task and $\alpha$ is responsible for the strength of the sparse regularization.
@@
$s_{i}=\sigma((\pi_{i}-\log(\log U_{1}/\log U_{2})/{\tau})\text{ with }U_{1},U_{2}\sim\text{U}(0,1),$ (2)
@@
$m_{i}=[\mathds{1}_{s_{i}>\gamma}-s_{i}]_{stop}+s_{i},$ (3)
where the threshold $\gamma$ is set to 0.5 by default, $\mathds{1}$ is the indicator function and $_{stop}$ is the stop gradient operator. Accordingly, the binary mask $m_{i}=1$, if the probability $s_{i}$ to keep this weight exceeds $0.5$ and $m_{i}=0$ otherwise. The final parameters of the resulting subnetwork are defined as $\mathbf{W^{\prime}=(w^{1}\odot m^{1},w^{2}\odot m^{2},...,w^{L}\odot m^{L})}$.
@@ 5 Our Approach
We hypothesize that instances with the same spurious attribute are nearby in representation space, and that spurious attributes induce clusters (cf. illustration in Fig.3 A). We verify this hypothesis for multiple attributes on the CelebA dataset (Sec. 5.1). Our approach centers around two key ideas: First, we identify clusters in the latent space and move instances with the same spurious attributes away from each other and samples of the same class closer to each other. Second, we reduce the representation capacity of the network to a subnetwork that is less prone to spurious correlations. Specifically, we learn a task-oriented subnetwork (Sec. 5.2) with a task-specific contrastive loss (Sec. 5.2.1) that optimizes the representation space. We describe the specific data selection and sampling procedure for negative and positive samples in the contrastive loss in Sec. 5.2.2. Finally, we present the overall training procedure in Sec. 5.3.
5.1 Analysis of ERM Representation Space for Multiple Spurious Attributes
@@
In the presence of spurious attributes, ERM tends to learn them as predictive features yielding decision boundaries that are aligned with the spurious feature (as exemplified on the two moons dataset in Fig. 1). We analyze the behavior of ERM in the presence of multiple spurious features on the CelebA dataset. We use an ERM model for predicting hair color (blond or not blond) having a training accuracy of 95%.
To analyze the feature space, we select three additional attributes (gender, age, presence of a beard) as potential spurious attributes. For a brief description of the CelebA dataset see Sec. 6.1 and Appendix A.1 for detailed information. Fig. 2 shows the t-SNE (Van der Maaten & Hinton, 2008) projection of a random subset of CelebA training data representations from the penultimate layer [^2] We discard samples with more than one of the selected attributes for easier interpretation.. The two classes are well separated in representation space (Fig. 2, left). Interestingly, there is a strong correlation between spurious attributes and classes in the representation space. Female correlates with blond hair, whereas the cluster for non-blond hair (orange cluster, left in Fig. 2) corresponds to young people and people with beard. Furthermore, within a class (non-blond hair), the two attributes partition the data into manifolds, even though the attribute labels are not used during training. Finally, unsupervised k-means clustering in representation space tends to cluster by attributes with high purity (Fig. 2, right). This means that when optimizing for class separation, instances from the same attribute group $a_{i}\in\mathcal{A}$ (e.g., persons with beard, young persons) lie close together in representation space, even though the group labels are not used during training.
@@
Tab. 1 shows the purity of classes (blond or non-blond hair) and spurious attributes (male or female) in clusters obtained from $k$-means clustering on representations of the CelebA training set. Each cluster is assigned a label based on the most frequent class/ spurious attribute. Purity is calculated as the number of correctly matched attribute and cluster labels divided by the total number of instances in that cluster. The table demonstrates that $k$-means clustering effectively groups samples by class label, but achieves even higher purity for spurious attributes. Fig. 2 (right) demonstrates $k$-means clustering ($k=8$) well defines clusters based on not only one but multiple spurious attributes. The critical case is when the model can potentially separate both, classes and attribute groups, i.e., when the model has high predictive performance and the latent space shows clear clusters for spurious attributes. Such a model will i) make predictions based on spurious attributes and ii) not generalize well to samples with attribute combinations that do not align with the spurious correlations.
Yang et al. (2024) investigate and provide a theoretical proof using a two-layer fully connected neural network trained on the CMNIST dataset (which has a spurious correlation between each class and a specific color) to mimic the early learning process of neural networks. The authors show that each group (defined by a label and its correlated spurious feature) can be separated early in training based on the model’s outputs. Our observations empirically show that even with a complex neural network, which is assumed to learn simple functions early in training and gradually build complexity over time, the representation space can still be separated by spurious features after training once spurious correlations are learned.
@@
Based on these observations, our goal is to ensure that spurious attributes do not define data manifolds and cannot be used as discriminative attributes.
5.2 Task-oriented Subnetworks
@@
To ensure that spurious attributes do not define data manifolds and cannot be used as discriminative attributes, we introduce a task-oriented subnetwork, and a contrastive loss that optimizes the representation space. More specifically, we extend the approach for extracting subnetworks (cf. Sec. 4) by incorporating a task-specific loss function to prevent the network from learning a grouping of samples based on spurious attributes. The loss for our task-oriented subnetwork is defined as
$\mathrm{L}=\mathrm{L_{mod}}+\beta\mathrm{L_{task}},$ (4)
@@
where $\mathrm{L_{mod}}$ is the modular loss (cf. Sec. 4), $\mathrm{L_{task}}$ the task-specific loss, and $\beta$ is a hyper-parameter to balance the loss terms.
@@
The choice of $\mathrm{L_{task}}$ in combination with a mask-training dataset $\mathcal{D}_{\text{task}}$ defines the specific task the subnetwork is optimized for. For example, Csordás et al. (2020) finds a subnetwork responsible for a particular class ($\mathcal{D}_{\text{task}}$ contains examples of this class, $\mathrm{L_{task}}$ = 0), and Park et al. (2023) identify a subnetwork that is more robust to a spurious feature ($\mathcal{D}_{\text{task}}$ contains instances misclassified by ERM, and $\mathrm{L_{task}}$ is a debiasing term). Details on these methods can be found in Appendix A.2.
@@
We detail our choice of $\mathrm{L_{task}}$ and $\mathcal{D}_{\text{task}}$ in Sec. 5.2.1 and 5.2.2, respectively. Fig. 3 illustrates the training procedure of our task-oriented subnetwork. Based on the assumption that instances of the same class (same color) lie in different clusters, and that spurious attributes induce clusters, we define a contrastive loss $\mathrm{L_{task}}$, which moves samples of the same class in different clusters closer together in representation space, and pushes samples of the same cluster apart.
@@
Figure 3: Assumption and overall idea. A. Instances from the same class (same colour) lie in different clusters apart in feature space. Instances with the spurious feature (solid border) are nearby in feature space, i.e., the spurious feature induces clusters and primarily defines the shape of the data manifold. B. Contrastive loss: For an anchor $x$, instances from the same class but different clusters constitute positive samples $x^{+}$, instances from the same cluster negative samples $x^{-}$. Positive samples are moved closer to the anchor, negative samples away from the anchor. In each iteration multiple samples act as anchors (with their positive and negative samples). C. Goal and effect of our contrastive loss. The spurious feature does not define data manifolds and cannot be used as discriminative feature between classes. Note: Our approach does not assume (and need) knowledge about the spurious feature(s).
@@ 5.2.1 Contrastive Learning
We use a variant of contrastive loss (Khosla et al., 2020) to mitigate the tendency of ERM to form clusters based on spurious attributes. Different from the original work, our contrastive batch is defined not only by the class labels, but also by the location of the samples in representation space. Fig. 3 outlines the idea.
For each contrastive batch we randomly sample a set of anchors and for each anchor a corresponding set of positive and negative samples. We define positive samples as samples of the same class as the anchor, but from a different cluster than the anchor. We define negative samples as samples from the same cluster (independent of their class label). For anchor $(x,y)$ belonging to cluster $\mathrm{c}\subset\mathcal{C}$, the positive set of size P is $\left\{(x^{+},y^{+})\mid(x^{+},y^{+})\in\mathcal{C}\setminus\mathrm{c}\land y^{+}=y\right\}$ and the negative set of size N is $\left\{(x^{-},y^{-})\mid(x^{-},y^{-})\in\mathrm{c}\setminus(x,y)\right\}$. The contrastive loss term defined with respect to the anchor $(x,y)$ is
$\mathrm{l_{con}}(x)=-\sum\limits_{i=1}^{P}\log\frac{\exp(z^{\top}z_{i}^{+}/\tau)}{\sum_{P}\exp(z^{\top}z_{p}^{+}/\tau)+\sum_{N}\exp(z^{\top}z_{n}^{-}/\tau)},$ (5)
where $\tau>0$ is a scalar temperature hyper-parameter, $z,z^{+},z^{-}$ are the representations obtained from the penultimate layer of the corresponding inputs $x$, $x^{+}$ and $x^{-}$ from anchor, positives and negatives respectively. We follow the finding of  (Khosla et al., 2020) that locating the summation outside the log improves performance over summation inside the log. This loss forces the representations of samples of the same class across clusters to become more similar and more similar to the anchor, while samples representing potentially spurious attributes are pushed away from the anchor.
@@ 5.2.2 Data
PruSC does not require prior knowledge of spurious attributes and is applicable in the presence of multiple spurious attributes. We do not assume that group annotations are available but automatically infer groups to construct the de-biasing dataset $\mathcal{D}_{\text{task}}$. Based on the assumption that the representation space of ERM has regions corresponding to spurious attributes (cf. Sec. 5.1, and Fig. 2), we use unsupervised clustering to construct the spurious groups for our dataset $\mathcal{D}_{\text{task}}$.
@@
Given an ERM model, we calculate the embeddings of the training data $\mathcal{D}$ and apply $k$-means clustering on the embeddings. Each sample in the training set $\mathcal{D}$ is assigned to a cluster $c$, $(x,y,c)\in\mathcal{D}=\mathcal{X}\times\mathcal{Y}\times\mathcal{C}$, with input domain $\mathcal{X}$, set of class labels $\mathcal{Y}$, and set of clusters $\mathcal{C}$. We define the number of classes in the classification task as $|\mathcal{Y}|=K$ and the number of clusters $|\mathcal{C}|=k$.
@@
We distinguish clusters based on their class purity. A cluster $\mathrm{c}$ is called $i$-dominant for class $i$ if at least 90% of its samples belong to class $i$, i.e., a cluster is $i$-dominant, if $\exists i\in\mathcal{Y}$ such that $\frac{|\{(x,y,c)\in\mathrm{c}\mid y=i\}|}{|\mathrm{c}|}\geq 0.9$. Otherwise, $\mathrm{c}$ is neutral. A cluster must fall into one of the $K+1$ categories that are either neutral or $i$-dominant for some $i\in\mathcal{Y}$ and multiple different clusters may be $i$-dominant for the same class $i$. $\mathrm{C}^{i}$ is the set of $i$-dominant clusters (of the same class $i$) and $\mathrm{C}_{N}$ is the set of neutral clusters.
Based on the observation in Section 5.1 that cluster purity is higher w.r.t. attributes than w.r.t. to classes, we derive that minority instances in an $i$-dominant cluster are typically hard cases. Hard cases contain a spurious attribute, but not the spurious correlation, i.e., are assigned to a class other than the spuriously correlated one. Accordingly, the set of hard cases $h_{i}$ obtained from the representation space contains samples with labels $i$ assigned to clusters dominated by a different class:
$h_{i}=\{(x,y,c)\mid c\in\mathrm{C}^{j\neq i}\land y=i\}$ (6)
@@
To construct a class-balanced dataset $\mathcal{D}_{\text{task}}$ for subnetwork training, we determine a fixed number $p$ of samples for each class. We draw $p$ samples for class $i$ as follows: First, we take all samples from the hard cases set $h_{i}$. Second, we draw the remaining samples for class $i$ uniformly from all $i$-dominant and neutral clusters. That is, from each cluster in $\mathrm{C}^{i}$ and $\mathrm{C}_{N}$, we draw an equal number of $\frac{p-|h_{i}|}{|\mathrm{C}^{i}|+|\mathrm{C}_{N}|}$ samples [^3] In the exceptional case $|h_{i}|+|\mathrm{C}|=0$ (no neutral cluster exists and all clusters are dominant for some class other than $i$) we only use $h_{i}$, which may result in less than $p$ samples. We never encountered such a situation in our experiments. having class label $i$. We obtain a subset of the training data $\mathcal{D}_{\text{task}}$ that is class balanced with $p$ samples for each class.
By sub-sampling $\mathcal{D}_{\text{task}}$, we intend to construct a subset from the training data without spurious correlations. In order to mine hard cases without prior knowledge of spurious correlations or access to ERM misclassifications from pre-training, we sample minority cases from $i$- dominant clusters. These instances are supposed to be anti-correlated. In addition, we draw remaining samples for class $i$ uniformly from all $i$-dominant and neutral clusters to construct a balanced dataset, i.e., without a spurious correlation. Drawing an equal number of samples per class results in a balancing, both along (spurious) features and classes. We also use that balanced dataset $\mathcal{D}_{\text{task}}$ for fine-tuning the extracted subnetwork after pruning, in order to avoid re-learning the spurious correlations. The ablation results in Sec. 7.1 confirm that fine-tuning with $\mathcal{D}_{\text{task}}$ is an effective performance improvement.